Solve( D^2+ DD'-6D'^2)z= cos(2x+y)

1
Expert's answer
2021-11-17T15:31:33-0500

"(D^2+DD'-6D'^2)z=cos(2x+y)"

The auxiliary equation is "m^2+m-6=0"

"m^2-2m+3m-6=0"

"m(m-2)+3(m-2)=0"

"(m+3)(m-2)=0"

Solved to get "m=2 , -3"

"\therefore" the C.F is f1(y+2x)+xf2(y-3x)

P.I ="\frac{1}{D^2+DD'-6D'^2}cos(2x+y)"

Replacing D2=-a2, DD'=-ab, D'2=-b2

From a=2, b=1

D2=-4, DD'=-2, D'2=-1

"P.I=\frac{1}{D^2+DD'-6D'^2}cos(2x+y)"

Incase of failure of f(D,D')=f(a,b)=0 for a=2, b=1 in cos(2x+y)

We therefore differentiate f(D,D') with respect to D and multiply f(x,y) by x

"P.I=\frac{1}{2D+D'}xcos(2x+y)"

Now replace D by a=2 and D' by b =1

"P.I=\frac{1}{2(2)+(1)}xcos(2x+y)"

"=\frac{x\ cos(2x+y)}{5}"

"z=C.F+P.I"

"\therefore z=f_1(y+2x)+xf_2(y-3x)+\frac{x\ cos(2x+y)}{5}"